# Large Errors in Kinetic Temperature Measurements Using Particle Tracking Velocimetry

### v1

@@ Abstract
We report on random errors in kinetic temperature measurements due to position uncertainty in particle tracking velocimetry. Using simulated data, we isolate the effects of spatial resolution and the frame rate on particle velocity calculations. A sample of particle velocities is generated from a Maxwellian distribution at a known temperature. Particle positions are assigned randomly and roughened to match a prescribed spatial resolution. Velocities are restored using the two-frame tracking method, and the kinetic temperature is calculated. By repeating this procedure, we obtain the discrepancy as a function of the spatial resolution and frame rate. Results show that, under typical experimental conditions, the restored kinetic temperature can deviate significantly from the true value, with spatial resolution playing a more dominant role than the frame rate.
@@ I Introduction
Direct access to microscopic particle properties is possible only in special cases. In simulations, molecular dynamics (MD) simulations provide complete access to individual particles properties. In experiments, systems such as granular materials and dusty plasmas allow individual particles to be resolved. In these situations, video microscopy can be used to measure particle positions frame by frame, from which trajectories, velocities, and even accelerations can be reconstructed. This approach of extracting particle velocities from video data is known as Particle Tracking Velocimetry (PTV) [1, 2].
@@ II Methodology
We do not use MD simulations for this purpose. MD introduces several complications that are undesirable here. First, a thermostat maintains the temperature only within fluctuations that depend on MD parameters, adding an uncertainty unrelated to PTV. Second, MD includes interparticle collisions and the resulting accelerations, which would reintroduce acceleration-related errors that we specifically wish to exclude. Third, velocity distribution functions in two-dimensional dusty plasmas are known to deviate from a Gaussian [cite], which complicates the kinetic-theory definition of temperature. Our method avoids these issues by prescribing the velocity distribution directly.
@@
Figure 3: Contour plots of the fractional error in the kinetic temperature measured by PTV, as a function of camera frame rate and spatial resolution. Results are shown for three true kinetic temperatures: (a) 10 eV, (b) 1.0 eV, and (c) 0.10 eV. The error arises solely from finite spatial resolution, modeled here as rounding particle positions to a pixel grid. Red crosses mark two operating points from our dusty plasma experiment, at $(\nu,R)=(99~\mathrm{fps},\,24.39~\mu\mathrm{m/px})$ and $(294~\mathrm{fps},\,30.69~\mu\mathrm{m/px})$. The contours show that the uncertainty increases with frame rate and decreases with coarser spatial resolution, and that the magnitude of the error depends strongly on the true kinetic temperature. Note the large uncertainties for typical frame rates and resolutions (center of each panel), especially in the low-temperature case (panel (c)).
@@
    This quantization introduces an effective position uncertainty of one pixel. Both frames are sampled on the same pixel grid, and the only loss of information comes from this rounding.
@@ III Discussion
For these calculations we used a sample size of $N=1000$, which is a common number of microspheres in a two-dimensional particle cloud [5, 6, 7, 8, 9]. The particle mass was $m=2.9\times 10^{-13}\,\mathrm{kg}$, corresponding to melamine–formaldehyde microspheres of diameter $7.14\,\mu\mathrm{m}$, which are also used in our laboratory.
@@
Our main results are shown in Fig. 3. Fig. 3 presents three contour plots showing the discrepancy in the measured kinetic temperature for true temperatures of 10 eV, 1.0 eV, and 0.10 eV. These values span the range commonly encountered in dusty plasma experiments. To illustrate how the results apply to an actual experiment, we mark two points corresponding to our laboratory conditions: $R=24.39\,\mu\mathrm{m/px}$ at $99$ fps, and $R=30.69\,\mu\mathrm{m/px}$ at $294$ fps.
Two general trends are visible in all three panels. First, for a fixed frame rate, the error decreases for cameras with coarser resolution (larger values of $R$). This follows directly from the fact that quantization of particle positions is the only source of error included in our model. For example, in panel (a), the uncertainty corresponding to our point at $294$ fps and $30.69\,\mu\mathrm{m/px}$ is about $26\%$, whereas for the same frame rate but the coarser resolution of $5\,\mu\mathrm{m/px}$ the uncertainty is below $10\%$.
@@ IV Summary
We found that the magnitude of this uncertainty is significant. Under typical conditions for dusty plasma experiments, the measured temperature can deviate from the true value by hundreds or thousands of percent. We confirmed previous findings that increasing the camera frame rate—often done to improve temporal resolution—can actually increase this error. This occurs because at high frame rates, the particle displacement between frames becomes comparable to the pixel size, making the quantization error dominant.

### v2

@@ Abstract
We report on random errors in kinetic temperature measurements due to finite spatial resolution in particle tracking velocimetry. Using simulated data, we isolate the error caused by finite spatial resolution from other sources of uncertainty, such as particle acceleration and particle mismatch. A sample of particle velocities is generated from a Maxwellian distribution at a prescribed kinetic temperature. Particle positions are assigned randomly and discretized to match a prescribed spatial resolution. Velocities are reconstructed using the two-frame tracking method, and the resulting kinetic temperature is calculated and compared to the true kinetic temperature. Results show that under typical experimental conditions, the uncertainty in particle positions propagates into large errors in the velocity distribution and the measured kinetic temperature. We find that this might introduce errors ranging from tens of percent at high kinetic temperatures ($\sim 10$ eV) to thousands of percent at low temperatures ($\sim 0.1$ eV).
@@ I Introduction
Direct access to microscopic particle properties is possible only in special cases. In simulations, molecular dynamics (MD) simulations provide complete access to individual particles properties. In experiments, systems such as granular materials and dusty plasmas allow individual particles to be resolved. In these situations, video microscopy can be used to measure particle positions frame by frame, from which trajectories, velocities, and even accelerations can be reconstructed. This approach of extracting particle velocities from video data is known as Particle Tracking Velocimetry (PTV) [1, 2, 3, 4].
@@
These findings apply generally to PTV measurements in any physical system, including granular materials [6, 7], colloids [1, 8], fluid dynamics [9, 10], and biophysics [11]. In this work, we specifically evaluate these errors for conditions typical of dusty plasma experiments [12, 13, 14, 15, 16, 17, 18, 19].
@@ II Methodology
We do not use MD simulations for this purpose. MD introduces several complications that are undesirable here. First, a thermostat maintains the temperature only within fluctuations that depend on MD parameters, adding an uncertainty unrelated to PTV. Second, MD includes interparticle collisions and the resulting accelerations, which would reintroduce acceleration-related errors that we specifically wish to exclude. Third, the particles systems like dusty plasmas are known to deviate from Gaussian statistics [20, 21, 22, 23, 24], which complicates the kinetic-theory definition of temperature. Our method avoids these issues by prescribing the velocity distribution directly.
@@
Figure 3: Contour plots of the fractional error in the kinetic temperature measured by PTV, as a function of camera frame rate and spatial resolution. Results are shown for three true kinetic temperatures: (a) 10 eV, (b) 1.0 eV, and (c) 0.10 eV. The error arises solely from finite spatial resolution. The red cross ($\times$) marks the operating point for our first experimental condition ($\nu=99~\mathrm{fps}$, $R=24.39~\mu\mathrm{m/px}$), and the red diamond ($\diamond$) marks the second condition ($\nu=294~\mathrm{fps}$, $R=30.69~\mu\mathrm{m/px}$). The white contour lines represent isolines of constant fractional error. The uncertainty increases with frame rate and decreases with coarser spatial resolution (larger $R$). Note the large uncertainties for typical frame rates and resolutions (center of each panel), especially in the low-temperature case (panel (c)).
@@
    This discretization introduces an effective position uncertainty of one pixel. Both frames are sampled on the same pixel grid, and the only loss of information comes from this rounding.
@@ III Discussion
For these calculations we used a sample size of $N=1000$, which is a common number of microspheres in a two-dimensional particle cloud [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. The particle mass was $m=2.9\times 10^{-13}\,\mathrm{kg}$, corresponding to melamine–formaldehyde microspheres of diameter $7.14\,\mu\mathrm{m}$, which are also used in our laboratory.
@@
Our main results are shown in Fig. 3. Fig. 3 presents three contour plots showing the discrepancy in the measured kinetic temperature for true temperatures of 10 eV, 1.0 eV, and 0.10 eV. These values span the range commonly encountered in dusty plasma experiments. To illustrate how the results apply to an actual experiment, we mark two points corresponding to our laboratory conditions. The red cross ($\times$) indicates operation at $R=24.39\,\mu\mathrm{m/px}$ and frame rate $99$ fps, while the red diamond ($\diamond$) indicates operation at $R=30.69\,\mu\mathrm{m/px}$ and frame rate $294$ fps.
Two general trends are visible in all three panels. First, for a fixed frame rate, the error decreases for cameras with coarser resolution (larger values of $R$). This follows directly from the fact that discretization of particle positions is the only source of error included in our model. For example, in panel (a), the uncertainty corresponding to our point at $294$ fps and $30.69\,\mu\mathrm{m/px}$ is about $26\%$, whereas for the same frame rate but the coarser resolution of $5\,\mu\mathrm{m/px}$ the uncertainty is below $10\%$.
@@ IV Summary
We found that the magnitude of this uncertainty is significant. Under typical conditions for dusty plasma experiments, the measured temperature can deviate from the true value by hundreds or thousands of percent. We confirmed previous findings [3] that increasing the camera frame rate—often done to improve temporal resolution—can actually increase this error. This occurs because at high frame rates, the particle displacement between frames becomes comparable to the pixel size, making the discretization error dominant.